### v2

# The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability

Abstract
We prove that there is no preferential voting method satisfying the Condorcet winner and loser criteria, positive involvement (if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose), and $n$-voter resolvability (if $x$ initially ties for winning, then $x$ can be made the unique winner by adding some set of up to $n$ voters). This impossibility theorem holds for any positive integer $n$. In a previous note, we proved an analogous result assuming an additional axiom of ordinal margin invariance, which we now show is unnecessary for an impossibility theorem, at least if the desired voting method is defined for five-candidate elections.
1 Introduction
In the theory of preferential voting, where voters submit rankings of the candidates, there is a fundamental tradeoff when selecting a voting method between ensuring desired variable-voter properties, ensuring desired variable-candidate properties, and ensuring that ties are not too frequent. In this note, we prove an impossibility theorem illustrating this tradeoff. Here the relevant variable-voter axiom is that of positive involvement (see Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose (see Definition 1.3). According to Pérez (2001, p. 605), this “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] This seems too strong, however, as a more plausible minimum would be bullet-vote positive involvement: adding a voter who ranks a would-be winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose. Another weaker condition is positive-negative involvement: adding a voter who ranks a winner $x$ uniquely first and a loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win. On the candidate side, although the famous Condorcet winner and loser criteria (Definition 1.1-2) are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp. lose) in every two-candidate sub-election, then $x$ should still win (resp. lose) in the full election—or else the election outcome might be seen as too contingent on which other losing candidates happened to stay in the race. [^2] As in the case of positive involvement, so too in the case of the Condorcet criteria—especially the Condorcet loser criterion—the motivation mentioned in the main text is not meant to be decisive. Finally, for a formalization of the idea that a voting method should not make ties too frequent, here we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a new voter (Definition 1.4). As Tideman (1986, p. 25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.” In fact, we will ultimately consider, for any positive integer $n$, the axiom of $n$-voter resolvability, whereby any tied winner can be made the unique winner by adding some set of up to $n$ new voters.
2 Setup
Given infinite sets $\mathcal{V}$ and $\mathcal{X}$ of possible voters and possible candidates, respectively, a profile $\mathbf{P}$ is a function assigning to each voter $i$ in some nonempty finite set $V(\mathbf{P})\subseteq\mathcal{V}$ a strict weak order $\mathbf{P}(i)$ on some nonempty finite set $X(\mathbf{P})\subseteq\mathcal{X}$ of candidates. We call $\mathbf{P}$ linear if for each $i\in V(\mathbf{P})$, $\mathbf{P}(i)$ is a linear order, so there are no ties in $i$’s ranking. All results to follow continue to hold if all profiles are linear. If $\mathbf{P}$ and $\mathbf{Q}$ are profiles with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{Q})$, then $\mathbf{P}+\mathbf{Q}$ is the profile with $V(\mathbf{P}+\mathbf{Q})=V(\mathbf{P})\cup V(\mathbf{Q})$ defined in the obvious way: if $i\in V(\mathbf{P})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{P}(i)$, and if $i\in V(\mathbf{Q})$, then $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{Q}(i)$.
Given $x,y\in X(\mathbf{P})$, the margin of $x$ vs. $y$ in $\mathbf{P}$ is the number of voters who rank $x$ above $y$ minus the number who rank $y$ above $x$: $\text{Margin}_{\mathbf{P}}(x,y)=|\{i\in V(\mathbf{P})\mid(x,y)\in\mathbf{P}(i)\}|-|\{i\in V(\mathbf{P})\mid(y,x)\in\mathbf{P}(i)\}|$. The margin graph $\mathcal{M}(\mathbf{P})$ of $\mathbf{P}$ is the weighted directed graph whose set of vertices is $X(\mathbf{P})$ with an edge from $x$ to $y$ when $\text{Margin}_{\mathbf{P}}(x,y)>0$, weighted by $\text{Margin}_{\mathbf{P}}(x,y)$.
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$. Note that this builds an assumption of “universal domain” into the definition of a voting method.
Definition 1.
Let $F$ be a voting method.
  1. 1.
    $F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(x,y)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet winner), then $F(\mathbf{P})=\{x\}$.
  2. 2.
    $F$ satisfies the Condorcet loser criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
  3. 3.
    $F$ satisfies positive involvement (Saari 1995) if for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{x})$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{x}$ consists of a single voter who ranks $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
  4. 4.
    $F$ satisfies resolvability (Tideman 1986) if for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ with $V(\mathbf{P})\cap V(\mathbf{S})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{S})$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. [^3] In Holliday 2024, we used a weaker version of this axiom in the impossibility theorem, since there the weaker version sufficed: if $|F(\mathbf{P})|>1$, then there is some single-voter profile $\mathbf{S}$ such $|F(\mathbf{P}+\mathbf{S})|=1$. But Tideman’s original axiom (here stated in a variable-voter setting) requires that if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. This version facilitates the proof in this paper. In Holliday 2024, we also discussed asymptotic resolvability, but without the assumption of ordinal margin invariance made in that paper, asymptotic resolvability is more difficult to work with, so we do not discuss it here.
Each of the axioms in Definition 1 is independent of the others. The Borda count (Borda 1781) satisfies all of the axioms except the Condorcet winner criterion. Minimax (Simpson 1969, Kramer 1977) satisfies all of them except the Condorcet loser criterion. [^4] Minimax does not satisfy resolvability if we restrict its domain to linear profiles (so the added voter must submit a linear order), but the related leximax method (see Brandt et al. 2025), wherein the winning candidates are those whose vectors of margins, sorted in increasing order, are lexicographically largest, does satisfy resolvability when its domain is restricted (or not restricted) to linear profiles, as well as the Condorcet winner criterion and positive involvement (for all profiles). Ranked Pairs (Tideman 1987) satisfies all of them except positive involvement. Split Cycle (Holliday and Pacuit 2023) satisfies all of them except resolvability.
Below we will prove that there is no voting method satisfying all of the axioms in Definition 1.555Among the corollaries of our main theorem is that there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement (Fishburn and Brams 1983, Pérez 2001) (if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\prime})$), resolvability, and neutral reversal (Saari 2003) (adding two voters with reversed linear orders does not change the set of winners); this uses the fact that if a voting method satisfies neutral reversal, then it satisfies positive involvement if and only if it satisfies negative involvement (see Ding et al. 2025, Prop. 3.18). The proof involves reasoning about five-candidate profiles. If we restrict the domain of voting methods to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax (Simpson 1969, Kramer 1977) voting method. For four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance (Holliday 2024), but we leave open whether they are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined for four-candidate profiles.
The impossibility proof to follow makes use of the following concept and lemma.
Definition 2.
Given a profile $\mathbf{P}$, the defensible set $D(\mathbf{P})$ of $\mathbf{P}$ is the set of candidates $x\in X(\mathbf{P})$ such that for all $y\in X(\mathbf{P})$, there is a $z\in X(\mathbf{P})$ such that $\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)$.
The following lemma is essentially Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof here for convenience.
Lemma 3 (Moulin 1988, Pérez 1995).
Let $F$ be a voting method satisfying positive involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$.
Proof.
Let $F$ be such a method, and suppose $x\in F(\mathbf{P})$. For contradiction, suppose $x\not\in D(\mathbf{P})$. Then there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, which implies $y\neq x$. Let $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, so $k<\text{Margin}_{\mathbf{P}}(y,x)$. If $k<0$, then $y$ is the Condorcet winner in $\mathbf{P}$, so $F(\mathbf{P})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P})$. Thus, $k\geq 0$. By our assumption about $\mathbf{P}$ in the lemma, $k+1<\text{Margin}_{\mathbf{P}}(y,x)$. Let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$. Hence by repeated application of positive involvement, $x\in F(\mathbf{P}^{\prime})$. But $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P}^{\prime})$. Thus, $x\in D(\mathbf{P})$.∎
3 Result
Theorem 4.
There is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and resolvability.
Proof.
Suppose for contradiction that there is such a voting method $F$. Let $\mathbf{P}_{1}$ be the profile in Figure 1, whose margin graph is $\mathcal{M}_{1}$ in Figure 2.666Rather than using this specific $\mathbf{P}_{1}$, one may reason more abstractly as follows. By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$. Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $10$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$. Then let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: e.g., $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 28 new voters who rank $a$ above $d$, $d$ above $b$, etc.; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $8$ voters who rank $d$ above $a$, $a$ above $e$, etc. Note there are indeed such $8$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $10$ voters in $\mathbf{P}_{1}$ who rank $d$ above $a$, $a$ above $c$, etc. Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
63 39 35 24 21 22 11 10 14 8 6 4
$a$ $b$ $e$ $b$ $c$ $e$ $e$ $d$ $d$ $d$ $b$ $a$
$c$ $a$ $b$ $a$ $e$ $d$ $d$ $a$ $c$ $a$ $d$ $e$
$d$ $c$ $d$ $c$ $b$ $b$ $c$ $c$ $e$ $e$ $a$ $c$
$e$ $e$ $c$ $d$ $d$ $a$ $b$ $b$ $b$ $c$ $c$ $b$
$b$ $d$ $a$ $e$ $a$ $c$ $a$ $e$ $a$ $b$ $e$ $d$
Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4
$\mathcal{M}_{1}$
aabbccddee87879595553311454551519999979777
+\;+ 28 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee595912312323233131272717177979717169692121 == 8 $d$ $a$ $e$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee676711511531313939191925257171636377771313
$\mathcal{M}_{3}$
aabbccddee676711511531313939191925257171636377771313
+\;+ 26 $b$ $d$ $e$ $c$ $a$ ==\;\; $\mathcal{M}_{4}$aabbccddee939389895757131377114545373751513939 == 10 $d$ $a$ $c$ $b$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1031037979676723231717993535474741412929
Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray
Now in $\mathbf{P}_{1}$, $a$ is the only defensible candidate who is not a Condorcet loser (see $\mathcal{M}_{1}$ in Figure 2). Thus, by the Condorcet loser criterion and Lemma 3, $F(\mathbf{P}_{1})=\{a\}$. Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement together imply $a\in F(\mathbf{P}_{2})$. Then by resolvability, there is some “empty” [^7] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc. The analogous convention applies to $\mathbf{S}_{4}$ below if $F(\mathbf{P}_{4}+\mathbf{S}_{2})=\{b\}$. or single-voter profile $\mathbf{S}_{2}$ such that $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$}. Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$. Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3. Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$. Then by resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$. Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$. Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3. Therefore, we have a contradiction.∎
The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
Theorem 5.
For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and $n$-voter resolvability.
Proof.
Replace each voter in the profiles $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of all margins in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
Thus, for any method satisfying the Condorcet winner and loser criteria and positive involvement, no matter the size of $n$, sometimes a tied winner cannot be made a unique winner even by the addition of up to $n$ voters.
Acknowledgements
I thank Emma Boniface and Diego Bejarano for helpful discussions of the problem posed at the end of Holliday 2024.