### v5

# The incompatibility of the Condorcet winner and loser criteria with positive involvement and resolvability

Abstract
We prove that there is no preferential voting method satisfying the Condorcet winner and loser criteria, positive involvement (if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose), and $n$-voter resolvability (if $x$ initially ties for winning, then $x$ can be made the unique winner by adding some set of up to $n$ voters). This impossibility theorem holds for any positive integer $n$. It also holds if either the Condorcet loser criterion is replaced by independence of clones or positive involvement is replaced by negative involvement.
1 Introduction
In the theory of preferential voting, where voters submit rankings of the candidates, there is a fundamental tradeoff when selecting a voting method between ensuring desired variable-voter properties, ensuring desired variable-candidate properties, and ensuring that ties are not too common. In this note, we prove an impossibility theorem illustrating this tradeoff. The relevant variable-voter axiom is that of positive involvement (Saari 1995, Pérez 2001, Holliday and Pacuit 2021): if a candidate $x$ wins in an initial preference profile, then adding a voter who ranks $x$ uniquely first cannot cause $x$ to lose. Alternatively, we can use negative involvement (Fishburn and Brams 1983, Pérez 2001): if a candidate $x$ loses in an initial preference profile, then adding a voter who ranks $x$ uniquely last cannot cause $x$ to win. According to Pérez (2001, p. 605), these “may be seen as the minimum to require concerning the coherence in the winning set when new voters are added.” [^1] However, a weaker requirement is positive-negative involvement: adding a voter who ranks an initial winner $x$ uniquely first and an initial loser $y$ uniquely last cannot cause $x$ to lose and $y$ to win. Another weaker requirement is bullet-vote positive involvement: adding a voter who ranks an initial winner $x$ uniquely first—and leaves all other candidates unranked or ranks them as tied after xx—cannot cause $x$ to lose. On the candidate side, although the famous Condorcet winner and loser criteria are not explicitly variable-candidate axioms, they can be motivated by considering candidates dropping out or joining the race: one might think that if $x$ would win (resp. lose) in every two-candidate sub-election, then $x$ should still win (resp. lose) in the full election—or else the election outcome might be seen as too contingent on which other losing candidates happened to stay in the race. Instead of the Condorcet loser criterion, we can use the explicitly variable-candidate axiom of independence of clones (Tideman 1987). Finally, for a formalization of the idea that a voting method should not make ties too common, we use the resolvability axiom that if $x$ is initially tied for winning, then $x$ can be made the unique winner by adding a single new voter. As Tideman (1986, p. 25) puts it, “Resolvability may be thought of as a condition that ties happen only “by coincidence”, every tie being within one vote of being broken.” In fact, we will ultimately consider, for any positive integer $n$, the axiom of $n$-voter resolvability, whereby any tied winner can be made the unique winner by adding some set of up to $n$ new voters.
2 Setup
Given infinite sets $\mathcal{V}$ and $\mathcal{X}$ of possible voters and possible candidates, respectively, a profile $\mathbf{P}$ is a function assigning to each voter $i$ in some nonempty finite set $V(\mathbf{P})\subseteq\mathcal{V}$ a strict weak order $\mathbf{P}(i)$ on some nonempty finite set $X(\mathbf{P})\subseteq\mathcal{X}$ of candidates. We call $\mathbf{P}$ linear if for each $i\in V(\mathbf{P})$, $\mathbf{P}(i)$ is a linear order. All results to follow continue to hold if all profiles are linear. If $\mathbf{P}$ and $\mathbf{Q}$ are profiles with $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{Q})$, then $\mathbf{P}+\mathbf{Q}$ is the profile with $V(\mathbf{P}+\mathbf{Q})=V(\mathbf{P})\cup V(\mathbf{Q})$ such that for $i\in V(\mathbf{P})$, $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{P}(i)$, and for $i\in V(\mathbf{Q})$, $(\mathbf{P}+\mathbf{Q})(i)=\mathbf{Q}(i)$.
Given $x,y\in X(\mathbf{P})$, the margin of $x$ vs. $y$ in $\mathbf{P}$ is the number of voters who rank $x$ above $y$ minus the number who rank $y$ above $x$: $\text{Margin}_{\mathbf{P}}(x,y)=|\{i\in V(\mathbf{P})\mid(x,y)\in\mathbf{P}(i)\}|-|\{i\in V(\mathbf{P})\mid(y,x)\in\mathbf{P}(i)\}|$. The margin graph $\mathcal{M}(\mathbf{P})$ of $\mathbf{P}$ is the weighted directed graph whose set of vertices is $X(\mathbf{P})$ with an edge from $x$ to $y$ when $\text{Margin}_{\mathbf{P}}(x,y)>0$, weighted by $\text{Margin}_{\mathbf{P}}(x,y)$.
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$. This builds an assumption of “universal domain” into the definition of a voting method.
Definition 1.
Let $F$ be a voting method.
  1. 1.
    $F$ satisfies the Condorcet winner criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(x,y)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet winner), then $F(\mathbf{P})=\{x\}$.
  2. 2.
    $F$ satisfies the Condorcet loser criterion (Condorcet 1785) if for every profile $\mathbf{P}$ and $x\in X(\mathbf{P})$, if $\text{Margin}_{\mathbf{P}}(y,x)>0$ for all $y\in X(\mathbf{P})\setminus\{x\}$ (so $x$ is a Condorcet loser), then $x\not\in F(\mathbf{P})$.
  3. 3.
    $F$ satisfies positive involvement (Saari 1995) if for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{x})$, if $x\in F(\mathbf{P})$ and $\mathbf{P}^{x}$ consists of a single voter who ranks $x$ uniquely first, then $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
  4. 4.
    $F$ satisfies resolvability (Tideman 1986) if for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ with $V(\mathbf{P})\cap V(\mathbf{S})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{S})$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. [^2] In Holliday 2024, we used a weaker version of this axiom in the impossibility theorem, since there the weaker version sufficed: if $|F(\mathbf{P})|>1$, then there is some single-voter profile $\mathbf{S}$ such $|F(\mathbf{P}+\mathbf{S})|=1$. But Tideman’s original axiom (here stated in a variable-voter setting) requires that if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is some single-voter profile $\mathbf{S}$ such that $F(\mathbf{P}+\mathbf{S})=\{x\}$. This version facilitates the proof in this paper. In Holliday 2024, we also discussed asymptotic resolvability, but without the assumption of ordinal margin invariance made in that paper, asymptotic resolvability is more difficult to work with, so we do not discuss it here.
Each of the axioms in Definition 1 is independent of the others. The Borda count (Borda 1781) satisfies all the axioms except the Condorcet winner criterion. Minimax (Simpson 1969, Kramer 1977) satisfies all of them except the Condorcet loser criterion. [^3] Minimax does not satisfy resolvability if we restrict its domain to linear profiles (so the added voter must submit a linear order), but the related leximax method (see Brandt et al. 2025)—wherein the winning candidates are those whose vectors of margins, sorted in increasing order, are lexicographically largest—does satisfy resolvability when its domain is restricted (or not restricted) to linear profiles, as well as the Condorcet winner criterion and positive involvement (for all profiles). Ranked Pairs (Tideman 1987) satisfies all of them except positive involvement. Split Cycle (Holliday and Pacuit 2023) satisfies all of them except resolvability.
Below we prove that there is no voting method satisfying all the axioms in Definition 1. The proof involves reasoning about profiles with five candidates. It extends the strategy of Holliday 2024, which showed that for four candidates, the axioms are jointly inconsistent with an additional axiom of ordinal margin invariance. We leave open whether the axioms are jointly satisfiable without ordinal margin invariance if we only ask for a voting method defined on four-candidate profiles. If we restrict the domain to two or three-candidate profiles, then the axioms are jointly satisfied by the Minimax voting method.
The impossibility proof in the next section makes use of the following concept (cf. Kasper et al. 2019).
Definition 2.
Given a profile $\mathbf{P}$, the defensible set $D(\mathbf{P})$ of $\mathbf{P}$ is the set of candidates $x\in X(\mathbf{P})$ such that for all $y\in X(\mathbf{P})$, there is a $z\in X(\mathbf{P})$ such that $\text{Margin}_{\mathbf{P}}(z,y)\geq\text{Margin}_{\mathbf{P}}(y,x)$.
The following lemma is essentially Lemma 3 of Pérez 1995 (cf. Lemma 1 of Pérez 2001), which is in turn an adaptation of a similar result of Moulin 1988, but we include a proof for convenience.
Lemma 3 (Moulin 1988, Pérez 1995).
Let $F$ be a voting method satisfying positive involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$.
Proof.
Let $F$ be such a method and $x\in F(\mathbf{P})$. For contradiction, suppose $x\not\in D(\mathbf{P})$. Then there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, which implies $y\neq x$. Let $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, so $k<\text{Margin}_{\mathbf{P}}(y,x)$. If $k<0$, then $y$ is the Condorcet winner in $\mathbf{P}$, so $F(\mathbf{P})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P})$. Thus, $k\geq 0$. By our assumption about $\mathbf{P}$ in the lemma, $k+1<\text{Margin}_{\mathbf{P}}(y,x)$. Let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by adding $k+1$ voters who rank $x$ uniquely first and $y$ uniquely second, followed by any linear order of $X(\mathbf{P})\setminus\{x,y\}$. Hence by repeated application of positive involvement, $x\in F(\mathbf{P}^{\prime})$. But $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, so $F(\mathbf{P}^{\prime})=\{y\}$ by the Condorcet winner criterion, which contradicts $x\in F(\mathbf{P}^{\prime})$. Thus, $x\in D(\mathbf{P})$.∎
3 Result
Theorem 4.
There is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and resolvability.
Proof.
Suppose for contradiction that there is such a voting method $F$. Let $\mathbf{P}_{1}$ be a profile as in Figure 1, whose margin graph is $\mathcal{M}_{1}$ in Figure 2.444Rather than using this specific $\mathbf{P}_{1}$, one may reason more abstractly as follows. By Debord’s Theorem (Debord 1987), there is a profile $\mathbf{Q}$ whose margin graph is $\mathcal{M}_{1}$. Now let $\mathbf{Q}^{+}$ be obtained from $\mathbf{Q}$ by adding, for each ranking of the five candidates, $7$ new voters with that ranking; then the margin graph of $\mathbf{Q}^{+}$ is still $\mathcal{M}_{1}$, and $\mathbf{Q}^{+}$ contains all the rankings needed for the rest of the argument, so one may use $\mathbf{Q}^{+}$ in place of $\mathbf{P}_{1}$. After the initial construction of this proof using another version of $\mathcal{M}_{1}$ and different numbers of rankings to add/subtract to obtain an appropriate sequence $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ of margin graphs for the proof, an integer linear program was used to find a small profile $\mathbf{P}_{1}$ and numbers of rankings to add/subtract to keep the total number of voters down while still satisfying all the constraints needed for each step of the proof. Even fewer voters suffice if, at the expense of elegance, we allow more distinct ranking types in the profiles. Let $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$ be obtained from $\mathbf{P}_{1}$ by adding and subtracting the profiles shown in the equations in Figure 2: $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding 26 new voters with the ranking adbecadbec; $\mathbf{P}_{3}$ is obtained from $\mathbf{P}_{2}$ by subtracting $7$ voters with the ranking dacebdaceb, etc. Note there are indeed such $7$ voters in $\mathbf{P}_{1}$, and for the purposes of $\mathbf{P}_{5}$, there are $7$ voters in $\mathbf{P}_{1}$ with the ranking dbacedbace. Then $\mathcal{M}_{2},\dots,\mathcal{M}_{5}$ are the margin graphs of $\mathbf{P}_{2},\dots,\mathbf{P}_{5}$, respectively.
69 64 46 20 18 2
$d$ $e$ $b$ $c$ $d$ $e$
$a$ $b$ $c$ $d$ $b$ $d$
$c$ $a$ $a$ $e$ $a$ $c$
$e$ $c$ $e$ $b$ $c$ $b$
$b$ $d$ $d$ $a$ $e$ $a$
Figure 1: the profile $\mathbf{P}_{1}$ used in the proof of Theorem 4. The number at the top of each column indicates the number of voters submitting the ranking in that column.
$\mathcal{M}_{1}$
aabbccddee8383114747818137371141418787919155
+\;+ 26 $a$ $d$ $b$ $e$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee109109272773735555636315156161252521216565 == 7 $d$ $a$ $c$ $e$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee102102343466666262707022225454181814145858
$\mathcal{M}_{3}$
aabbccddee102102343466666262707022225454181814145858
+\;+ 23 $b$ $d$ $e$ $a$ $c$ ==\;\; $\mathcal{M}_{4}$aabbccddee12512511114343858593935531311137373535 == 7 $d$ $b$ $a$ $c$ $e$ ++\;\; $\mathcal{M}_{5}$aabbccddee1181181818363678788686121266242430304242
Figure 2: margin graphs used in the proof of Theorem 4 with defensible candidates shaded gray. An edge from candidate $x$ to candidate $y$ labeled by $k$ indicates that the margin of $x$ vs. $y$ is $k$.
Now in $\mathbf{P}_{1}$, $a$ is the only defensible candidate who is not a Condorcet loser (see $\mathcal{M}_{1}$ in Figure 2). Thus, by the Condorcet loser criterion and Lemma 3, $F(\mathbf{P}_{1})=\{a\}$. Since $\mathbf{P}_{2}$ is obtained from $\mathbf{P}_{1}$ by adding voters all of whom rank $a$ uniquely first, $F(\mathbf{P}_{1})=\{a\}$ and repeated application of positive involvement together imply $a\in F(\mathbf{P}_{2})$. Then by resolvability, there is some “empty” [^5] Here $\mathbf{S}_{2}$ is not literally an empty profile; if we already have $F(\mathbf{P}_{2})=\{a\}$, then let $\mathbf{P}_{2}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{2}$, $\mathbf{P}_{3}+\mathbf{S}_{2}$’ denote $\mathbf{P}_{3}$, etc. The analogous convention applies to $\mathbf{S}_{4}$ below if $F(\mathbf{P}_{4}+\mathbf{S}_{2})=\{b\}$. or single-voter profile $\mathbf{S}_{2}$ such that $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$}. Since $\mathbf{P}_{2}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{2}+\mathbf{S}_{2})=\{a\}$} and positive involvement together imply $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$. Since any difference between weights of two distinct edges in $\mathcal{M}_{3}$ is of size at least $4$, no matter which single-voter (or empty) profile $\mathbf{S}_{2}$ is, $b$ is the only defensible candidate besides $d$ in $\mathbf{P}_{3}+\mathbf{S}_{2}$, and $\mathbf{P}_{3}+\mathbf{S}_{2}$ satisfies the margin separation assumption of Lemma 3, so $d\not\in F(\mathbf{P}_{3}+\mathbf{S}_{2})$ implies $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ by Lemma 3. Then since $\mathbf{P}_{4}+\mathbf{S}_{2}$ is obtained from $\mathbf{P}_{3}+\mathbf{S}_{2}$ by adding voters all of whom rank $b$ uniquely first, $F(\mathbf{P}_{3}+\mathbf{S}_{2})=\{b\}$ and positive involvement together imply $b\in F(\mathbf{P}_{4}+\mathbf{S}_{2})$. Then by resolvability, there is some “empty” or single-voter profile $\mathbf{S}_{4}$ such that $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$. Since $\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4}$ is obtained from $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ by adding voters all of whom rank $d$ uniquely first, $F(\mathbf{P}_{4}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{b\}$ and positive involvement together imply $d\not\in F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})$. Finally, since any difference between weights of two distinct edges in $\mathcal{M}_{5}$ is of size at least $6$, no matter which single-voter (or empty) profiles $\mathbf{S}_{2}$ and $\mathbf{S}_{4}$ are, $d$ is the only defensible candidate in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, and $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$ satisfies the margin separation assumption of Lemma 3, so $F(\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4})=\{d\}$ by Lemma 3. Therefore, we have a contradiction.∎
Remark 5.
Inspection of the proof of Theorem 4 shows that resolvability can be replaced by quasi-resoluteness (in the terminology of Holliday and Pacuit 2023): $|F(\mathbf{P})|=1$ for any profile $\mathbf{P}$ that is uniquely weighted in the sense that for all $x,y,x^{\prime},y^{\prime}\in X(\mathbf{P})$ with $x\neq y$ and $x^{\prime}\neq y^{\prime}$, if $(x,y)\neq(x^{\prime},y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,y)\neq\text{Margin}_{\mathbf{P}}(x^{\prime},y^{\prime})$. It also shows that the package of positive involvement and resolvability can be replaced by strict positive involvement (in the terminology of Dominik Peters, personal communication), which is defined like positive involvement but with $F(\mathbf{P}+\mathbf{P}^{x})=\{x\}$ instead of $x\in F(\mathbf{P}+\mathbf{P}^{x})$.
The same proof strategy can be used to show that for each positive integer $n$, resolvability can be replaced by $n$-voter resolvability: for every profile $\mathbf{P}$, if $|F(\mathbf{P})|>1$, then for each $x\in F(\mathbf{P})$, there is a profile $\mathbf{Q}$ with $|V(\mathbf{Q})|\leq n$, $V(\mathbf{P})\cap V(\mathbf{Q})=\varnothing$, and $X(\mathbf{P})=X(\mathbf{Q})$ such that $F(\mathbf{P}+\mathbf{Q})=\{x\}$.
Theorem 6.
For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, positive involvement, and $n$-voter resolvability.
Proof.
Replace each voter in $\mathbf{P}_{1},\dots,\mathbf{P}_{5}$ by $n$ copies of that voter, which results in a multiplication of each margin in $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ by $n$, and then the proof from Theorem 4 goes through as before but now with $|V(\mathbf{S}_{2})|\leq n$ and $|V(\mathbf{S}_{4})|\leq n$.∎
These proofs can also be adapted to use negative involvement (Fishburn and Brams 1983) instead of positive involvement: for every profile $\mathbf{P}$, $x\in X(\mathbf{P})$, and profile $\mathbf{P}^{\mathord{\downarrow}x}$ with $V(\mathbf{P})\cap V(\mathbf{P}^{\mathord{\downarrow}x})=\varnothing$ and $X(\mathbf{P})=X(\mathbf{P}^{\mathord{\downarrow}x})$, if $x\not\in F(\mathbf{P})$ and $\mathbf{P}^{\mathord{\downarrow}x}$ consists of a single voter who ranks $x$ uniquely last, then $x\not\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$. Contrapositively, if $x\in F(\mathbf{P}+\mathbf{P}^{\mathord{\downarrow}x})$, then $x\in F(\mathbf{P})$.
Theorem 7.
There is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and resolvability.
Proof.
A variant of Lemma 3 holds for negative involvement (cf. Pérez 2001, Lemma 1): Let $F$ be a voting method satisfying negative involvement and the Condorcet winner criterion. Let $\mathbf{P}$ be a profile in which for all $x,x^{\prime},y,y^{\prime}\in X(\mathbf{P})$ with $x\neq x^{\prime}$ and $y\neq y^{\prime}$, if $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})$, then $\text{Margin}_{\mathbf{P}}(x,x^{\prime})>\text{Margin}_{\mathbf{P}}(y,y^{\prime})+1$. Additionally assume that for each $x\in X(\mathbf{P})\setminus D(\mathbf{P})$, there is a $y\in X(\mathbf{P})$ such that for every $z\in X(\mathbf{P})$, $\text{Margin}_{\mathbf{P}}(z,y)<\text{Margin}_{\mathbf{P}}(y,x)$, and where $k=\max\{\text{Margin}_{\mathbf{P}}(z,y)\mid z\in X(\mathbf{P}),z\neq y\}$, there are at least $k+1$ voters whose rankings have $y$ uniquely second-to-last, followed by $x$ uniquely last. Then $F(\mathbf{P})\subseteq D(\mathbf{P})$. The proof is the same as that of Lemma 3 except we obtain $\mathbf{P}^{\prime}$ from $\mathbf{P}$ by removing $k+1$ voters of the kind described above. Then since the proof of Lemma 3 begins with $x\in F(\mathbf{P})$, repeated application of negative involvement in its contrapositive form implies $x\in F(\mathbf{P}^{\prime})$. Removing a ranking $L$ has the same effect on margins as adding the reversed ranking $L^{-1}$, so as in the proof of Lemma 3, $y$ is the Condorcet winner in $\mathbf{P}^{\prime}$, yielding a contradiction.
Now we adapt the proof of Theorem 4 by adding 147 copies of each linear order of $X(\mathbf{P}_{1})$ to $\mathbf{P}_{1}$, which ensures that the variant of Lemma 3 is applicable when needed in the proof. [^6] By starting with at least $147$ copies of each linear order in $\mathbf{P}_{1}$, we ensure that there are at least $121$ copies of each linear order in $\mathbf{P}_{3}+\mathbf{S}_{2}$ and in $\mathbf{P}_{5}+\mathbf{S}_{2}+\mathbf{S}_{4}$, which is greater than the maximal margins in those profiles, so the variant of Lemma 3 is applicable to those profiles. Of course, one can use an integer linear program to synthesize a new starting profile with far fewer voters such that all constraints needed for the proof still hold. Then instead of adding 26 voters with the ranking adbecadbec to obtain $\mathbf{P}_{2}$, we remove 26 voters with cebdacebda; instead of removing 7 voters with dacebdaceb from $\mathbf{P}_{2}$ to obtain $\mathbf{P}_{3}$, we add 7 voters with becadbecad; and so on for $\mathbf{P}_{4}$ (removing 23 voters with caedbcaedb) and $\mathbf{P}_{5}$ (adding 7 voters with ecabdecabd). The margins of $\mathcal{M}_{1},\dots,\mathcal{M}_{5}$ remain as in the proof of Theorem 4, and then the rest of the proof is the same except that we (i) invoke the variant of Lemma 3 for steps involving defensibility and (ii) use negative involvement instead of positive involvement.∎
By a scaling argument exactly analogous to that in the proof of Theorem 6, we can extend the proof of Theorem 7 to establish the following.
Theorem 8.
For each positive integer $n$, there is no voting method satisfying the Condorcet winner and loser criteria, negative involvement, and $n$-voter resolvability.
Finally, in response to a previous version of this note with the results above, Dominik Peters (personal communication) suggested that analogous impossibilities might hold with independence of clones (Tideman 1987) instead of the Condorcet loser criterion. Indeed, we can confirm that such a swap in Theorems 4 and 6 is possible, as well as in Theorems 78 if we add the axiom of block preservation (Holliday and Pacuit 2025): if the profile $\mathbf{L}$ consists of exactly one copy of each linear order of $X(\mathbf{P})$, then $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{L})$.
Recall that a set $C\subsetneq X(\mathbf{P})$ with $|C|\geq 2$ is a set of clones in $\mathbf{P}$ if for each $i\in V(\mathbf{P})$ and $x\in X(\mathbf{P})\setminus C$, either $i$ ranks $x$ above each $c\in C$ or $i$ ranks $x$ below each $c\in C$. Given a profile $\mathbf{P}$ and $c\in X(\mathbf{P})$, let $\mathbf{P}_{-c}$ be the profile with $X(\mathbf{P}_{-c})=X(\mathbf{P})\setminus\{c\}$ and $V(\mathbf{P}_{-c})=V(\mathbf{P})$ where for each $i\in V(\mathbf{P})$, $\mathbf{P}_{-c}(i)$ is the restriction of $\mathbf{P}(i)$ to $X(\mathbf{P}_{-c})$. Then $F$ satisfies independence of clones if for any profile $\mathbf{P}$, set $C$ of clones, $c\in C$, and $x\in X(\mathbf{P})\setminus C$, (i) $x\in F(\mathbf{P})$ iff $x\in F(\mathbf{P}_{-c})$, and (ii) $C\cap F(\mathbf{P})\neq\varnothing$ iff $C\cap F(\mathbf{P}_{-c})\neq\varnothing$.
Theorem 9.
For each positive integer $n$, there is no voting method satisfying the Condorcet winner criterion, independence of clones, positive involvement, and $n$-voter resolvability.
The same holds with positive involvement replaced by negative involvement plus block preservation.
Proof.
We prove the $n=1$ case. Reasoning as in the proof of Theorem 6 handles $n>1$. For positive involvement, consider a $\mathbf{Q}_{1}$ as in Figure 3, wherein $\{a,b,c,e\}$ is a set of clones. By Lemma 3, $F(\mathbf{Q}_{1})\subseteq\{a,d\}$. By the Condorcet winner criterion, $F(\mathbf{Q}_{1-b})=\{a\}$, so by independence of clones (part (i) or (ii)) and $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, we have $a\in F(\mathbf{Q}_{1})$. Then obtain $\mathbf{Q}_{2},\dots,\mathbf{Q}_{5}$ from $\mathbf{Q}_{1}$ by adding and subtracting the profiles in Figure 4. The rest of the proof is just like that of Theorem 4 but with $\mathbf{Q}_{i}$ instead of $\mathbf{P}_{i}$. For negative involvement, by block preservation, we can add as many blocks of all linear orders as we wish to $\mathbf{Q}_{1}$ to obtain a $\mathbf{Q}_{1}^{+}$ with $F(\mathbf{Q}_{1})\subseteq F(\mathbf{Q}_{1}^{+})\subseteq\{a,d\}$ by the variant of Lemma 3 in the proof of Theorem 7. Hence $F(\mathbf{Q}_{1})\subseteq\{a,d\}$, which implies $a\in F(\mathbf{Q}_{1})$ as above and therefore $a\in F(\mathbf{Q}_{1}^{+})$. Then starting from $\mathbf{Q}_{1}^{+}$, we reason with negative involvement instead of positive involvement as in the proof of Theorem 7.∎
62 60 42 23 19 3
$d$ $c$ $d$ $a$ $e$ $c$
$b$ $b$ $e$ $e$ $c$ $e$
$a$ $a$ $a$ $c$ $b$ $b$
$e$ $e$ $c$ $b$ $a$ $a$
$c$ $d$ $b$ $d$ $d$ $d$
Figure 3: the profile $\mathbf{Q}_{1}$ used in the proof of Theorem 9.
$\mathcal{M}_{1}$
aabbccddee45451181817979113535858511838311
+\;+ 16 7 2 $a$ $a$ $a$ $d$ $e$ $e$ $b$ $d$ $b$ $c$ $b$ $d$ $e$ $c$ $c$ ==\;\; $\mathcal{M}_{2}$aabbccddee7070262610610654544242606020202424667676 == 5 $d$ $e$ $a$ $c$ $b$ ++\;\; $\mathcal{M}_{3}$aabbccddee6565313111111159594747555515151919117171
$\mathcal{M}_{3}$
aabbccddee6565313111111159594747555515151919117171
+\,\;\;\;\,+\;\;\; 20 $b$ $d$ $c$ $a$ $e$ =\;\;\;=\;\;\;\;\;\,\, $\mathcal{M}_{4}$aabbccddee4545111113113179795567673535393921215151 == 11 $d$ $b$ $a$ $e$ $c$ ++\;\; $\mathcal{M}_{5}$aabbccddee343422221201206868161656564646282810104040
Figure 4: margin graphs used in the proof of Theorem 9 with defensible candidates shaded gray.
Acknowledgements
I thank Diego Bejarano and Emma Boniface for helpful discussions of the problem posed at the end of Holliday 2024 and Dominik Peters for helpful feedback on this paper.